# BUFFER-X: Towards Zero-Shot Point Cloud Registration in Diverse Scenes

### v1

@@ 4.3 Hierarchical inlier search
Intra-scale matching. First, we perform nearest neighbor-based mutual matching [47] between $\mathcal{F}^{\mathcal{P}}_{\xi}$ and $\mathcal{F}^{\mathcal{Q}}_{\xi}$, yielding matched correspondences $\mathcal{A}_{\xi}$ at each scale. Using $\mathcal{A}_{\xi}$, we extract the corresponding elements from $\mathcal{C}^{\mathcal{P}}_{\xi}$ and $\mathcal{C}^{\mathcal{Q}}_{\xi}$, denoted as $\widehat{\mathcal{C}}^{\mathcal{P}}_{\xi}$ and $\widehat{\mathcal{C}}^{\mathcal{Q}}_{\xi}$, and the sampled keypoints from ${\mathcal{P}_{\xi}}$ and ${\mathcal{Q}_{\xi}}$ as $\widehat{\mathcal{P}}_{\xi}$ and $\widehat{\mathcal{Q}}_{\xi}$, respectively (i.e., $|\mathcal{A}_{\xi}|=|\widehat{\mathcal{C}}^{\mathcal{P}}_{\xi}|=|\widehat{\mathcal{C}}^{\mathcal{Q}}_{\xi}|=|\widehat{\mathcal{P}}_{\xi}|=|\widehat{\mathcal{Q}}_{\xi}|$).
@@
[…]
@@
where ${\bm{n}}^{{\bm{p}}}={\bm{v}}^{{\bm{p}}}_{3}\times{\bm{z}}$, $\theta^{{\bm{p}}}=\cos^{-1}\left({\bm{v}}^{{\bm{p}}}_{3}\cdot{\bm{z}}\right)$, and $[\cdot]_{\times}$ denotes the skew operator (resp. ${\bm{R}}^{\bm{q}}$). Thus, once the yaw rotation between the two patches ${\bm{R}}_{\text{yaw}}$ is determined, the full 3D rotation can be obtained as ${\bm{R}}=\left({\bm{R}}^{\bm{q}}\right)^{\intercal}{\bm{R}}_{\text{yaw}}{\bm{R}}^{\bm{p}}$.
As explained by Ao  et al. [5], ${\bm{c}}^{{\bm{p}}}$ and ${\bm{c}}^{{\bm{q}}}$ follow discretized SO(2)-equivariant representation; thus, by finding the yaw rotation that maximizes circular cross-correlation between ${\bm{c}}^{{\bm{p}}}$ and ${\bm{c}}^{{\bm{q}}}$, we can estimate the relative SO(2) rotation ${\bm{R}}_{\text{yaw}}$. To this end, a 4D matching cost volume ${\bm{V}}\in\mathbb{R}^{H\times W\times W\times D}$ is constructed to represent the sector-wise differences between ${\bm{c}}^{{\bm{p}}}$ and ${\bm{c}}^{{\bm{q}}}$. Then, ${\bm{V}}$ is processed by a 3D cylindrical convolutional network (3DCCN) [6], mapping ${\bm{V}}$ to a score vector ${\bm{\beta}}$ of size $W$.
@@
Cross-scale consensus maximization. Then, using per-pair $({\bm{R}},{\bm{t}})$ estimates from all scales, the 3D point pairs with the largest cardinality across scales should be selected as the final inlier correspondences $\mathcal{I}$, ensuring cross-scale consistency. To achieve this, we formulate the cross-scale inlier selection as consensus maximization problem [68, 89].
@@
[…]
@@
Finally, $\mathcal{I}$ is given as input to a solver, such as RANSAC [27] or TEASER++ [77] to estimate $\hat{{\bm{R}}}$ and $\hat{{\bm{t}}}$. For a fair comparison with existing approaches, we use RANSAC.
@@ 4.4 Loss function and training
Loss functions. Unlike BUFFER, which was trained in four stages, our network follows a relatively simpler two-stage training process thanks to its detector-free nature. First, we train the feature discriminability of Mini-SpinNet descriptors using contrastive learning [82], followed by training $d$ in 7 to improve transformation estimation accuracy.
In particular, we employ the Huber loss [92] $\rho_{\text{Huber}}(\cdot)$ for training $d$ to remain robust to outliers[9], while balancing sensitivity to small errors, which is formulated as follows:
@@
Patch distribution augmentation. Furthermore, we propose an inter-patch point distribution augmentation to allow Mini-SpinNet to experience a wider variety of patch distribution patterns. Specifically, we empirically sample the radius within $[\frac{2}{3}r,\frac{4}{3}r]$ based on a uniform probability. As mentioned in Sec. 4.2, since $N_{\text{patch}}$ points within the radius are randomly selected as an input, a diverse set of patterns can be provided as $r$ varies.
@@ 5 Experiments
Datasets. As presented in Table 1, we designed our generalizability benchmark using eleven different datasets [88, 32, 81, 60, 69, 28, 56, 34, 71, 61] to ensure balanced consideration of the following aspects: a) variation in environmental scales (i.e., indoor and outdoor environments), b) different scanning patterns with different sensor types; see “Sensor Type” column in Table 1, c) acquisition setups, and d) diversity of geographic and cultural environments as the data was collected across Europe, Asia, and the USA (i.e., Oxford, KAIST, and MIT campuses, respectively). More details can be found in Appendices B and C.
Dataset Name Env. Sensor Type Acquisition Range [m] # Pairs
Indoor 3DMatch [88] Room RGB-D cameras Handheld 3.5 1,623
3DLoMatch [32] Room RGB-D cameras Handheld 3.5 1,781
[…]
ScanNet++F [81] Room FARO (Stationary LiDAR) Tripod 7.0 2,016
TIERS [60]
Room,
Interior
VLP16, OS1-64, OS0-128 Cart 110 870
[…]
KITTI [28] City HDL-64E Vehicle 80 555
ETH [56] Forest Rotating 2D Scanner Tripod 30 713
KAIST
(from HeLiPR [34])
Campus
[…]
Vehicle 240 1,991
MIT
(from Kimera-Multi [14])
Campus VLP16 Wheeled 100 230
Oxford
(from NewerCollege [61])
Campus OS1-64
Quadruped /
Handheld
100 301
Table 1: Overview of datasets used in our experiments, categorized by the environment, measurement sensor type, acquisition setup, average maximum range, and a number of test pairs. The superscript ✽ highlights solid-state LiDAR sensors to especially emphasize that our evaluation is not limited to conventional omnidirectional spinning LiDAR sensors, but contains different scanning patterns.
Training settings. Then, we only train the network on a single dataset, such as 3DMatch [88] or KITTI [28]. Using the same hyperparameters of BUFFER [5], we conducted a two-stage optimization (i.e., Mini-SpinNet is first trained, followed by training the 3DCCN) and we used Adam optimizer [35] with a learning rate of 0.001, a weight decay of 1e-6, and a learning rate decay of 0.5. We used NVIDIA GeForce RTX 3090 with AMD EPYC 7763 64-Core.
Testing settings. In the case of the existing datasets [88, 32, 69, 28, 56], we follow the conventional given pairs. The description of the newly employed datasets for evaluation can be found in Appendix B.
Evaluation Metrics. As a key metric, we use the success rate, which directly assesses the robustness of global registration [40]. Specifically, a registration is deemed successful if the translation and rotation errors are within $\tau_{\text{trans}}$ and $\tau_{\text{rot}}$, respectively [82]. For successful cases, we evaluated the performance using relative translation error (RTE) and relative rotation error (RRE), which are defined as follows:
@@
Conven- tional FPFH [63] + FGR [94] + [Uncaptioned image] 62.53 15.42 77.68 92.31 80.60 98.74 100.00 89.80 74.78 91.87 99.00 9.55
FPFH [63] + Quatro [43] + [Uncaptioned image] 8.22 1.74 9.88 97.27 86.57 99.10 100.00 91.46 79.57 51.05 91.03 10.73
FPFH [63] + TEASER++ [77] + [Uncaptioned image] 52.00 13.25 66.15 97.22 73.13 98.92 100.00 89.20 71.30 93.69 99.34 10.00
Deep learning-based FCGF [22] 88.18 40.09 72.90 88.69 55.96 0.00 0.00 0.00 0.00 54.98 0.00 15.00
@@
Ours 95.58 74.18 94.99 99.90 93.45 99.82 100.00 99.15 97.30 99.72 99.67 1.55
Table 2: Quantitative performance comparison on rate to evaluate generalization capability. Deep learning-based models were trained only on 3DMatch [88] and RANSAC was used with the maximum iteration of 50K. The icons represent oracle tuning ([Uncaptioned image]) to set proper voxel size and search radius to each dataset, and scale alignment ([Uncaptioned image]) to normalize dataset scales to be similar to that of 3DMatch data (e.g., the scale of KITTI, which typically uses a voxel size of 0.3 m, is adjusted to match the scale of 3DMatch, which typically uses a voxel size of 0.025 m, by dividing by $\frac{0.3}{0.025}$).
@@ 5.1 Analyses on the generalization
First, we demonstrate that existing methods struggle in achieving out-of-the-box generalization, leading to performance degradation due to the issues explained in Sec. 3. In this experiment, we mainly used renowned learning-based approaches: FCGF [22], Predator [32], GeoTransformer (GeoT for brevity) [59], BUFFER [5], and PARENet [80].
As shown in Table 2, models trained with small voxel sizes and search radii for indoor datasets exhibited substantial performance degradation in outdoor scenarios. In particular, as explained in Sec. 3.2.1, we observed that some approaches even did not work due to out-of-memory issues (i.e., N/A in Table 2) caused by an excessively large number of input points.
@@ 5.2 Performance comparison
Fig. 5: Relative translation error (RTE) and relative rotation error (RRE) of our approach to state-of-the-art methods, all trained on 3DMatch and tested on KITTI, with oracle tuning and scale alignment, corresponding to those in Table 2 under the + Refer to caption + Refer to caption setting. The **** annotations indicate measurements with a $p$-value $<10^{-4}$ after a paired $t$-test.
@@
Conventional. G-ICP [66] 8.56 0.22 37.95
FPFH [63] + FGR [94] 18.75 0.38 98.74
FPFH [63] + Quatro [43] 18.56 0.93 99.10
FPFH [63] + TEASER++ [77] 15.35 0.68 98.92
KISS-Matcher [41] 21.33 0.96 99.46
Learning-based 3DFeat-Net [82] 25.90 0.57 95.97
FCGF [22] 6.47 0.23 98.92
DIP [53] 8.69 0.44 97.30
Predator [32] 5.60 0.24 99.82
SpinNet [6] 9.88 0.47 99.10
CoFiNet [85] 8.20 0.41 99.82
D3Feat [8] 11.00 0.24 99.82
GeDi [54] 7.55 0.33 99.82
GeoTransformer [59] 7.40 0.27 99.82
BUFFER [5] 7.46 0.26 99.64
PARE-Net [80] 4.90 0.23 99.82
Proposed 7.74 0.27 99.82
Table 3: In-domain quantitative results in terms of relative translation error (RTE), relative rotation error (RRE), and success rate on KITTI [82].
@@ 5.3 Ablation study
Fig. 6: Success rate changes of state-of-the-art approaches based on the voxel size and search radius, where ★ indicates the best performance after tuning. (a) Impact of voxel size on success rate for models trained on 3DMatch with scale alignment, evaluated on KITTI. (b) Comparison of our BUFFER-X and BUFFER with varying search radius, highlighting that the optimal radius might differ even within the same dataset. For a fair comparison, BUFFER-X is evaluated in single-scale mode.
Learning-based detector vs. Farthest point sampling. Interestingly, FPS rather showed better performance than the learning-based keypoint detector in BUFFER[5], even in its training domain (i.e., in 3DMatch and 3DLoMatch). In particular, the performance gap sometimes becomes more pronounced in the out-of-domain scenes, demonstrating that robust cross-domain generalization can be achieved without the need for additional learning-based keypoint selection.
Fig. 7: Comparison of success rates between random sampling, learning-based keypoint detection in BUFFER [5], and our farthest point sampling (FPS) strategy, showing that FPS performs comparably or even better across various datasets.

### v3

@@ 4.3 Hierarchical inlier search
Intra-scale matching. First, we perform nearest neighbor-based mutual matching [Lowe04ijcv] between $\mathcal{F}^{\mathcal{P}}_{\xi}$ and $\mathcal{F}^{\mathcal{Q}}_{\xi}$, yielding matched correspondences $\mathcal{A}_{\xi}$ at each scale. Using $\mathcal{A}_{\xi}$, we extract the corresponding elements from $\mathcal{C}^{\mathcal{P}}_{\xi}$ and $\mathcal{C}^{\mathcal{Q}}_{\xi}$, denoted as $\widehat{\mathcal{C}}^{\mathcal{P}}_{\xi}$ and $\widehat{\mathcal{C}}^{\mathcal{Q}}_{\xi}$, and the sampled keypoints from $\mathcal{P}_{\xi}$} and $\mathcal{Q}_{\xi}$} as $\widehat{\mathcal{P}}_{\xi}$ and $\widehat{\mathcal{Q}}_{\xi}$, respectively (i.e., $|\mathcal{A}_{\xi}|=|\widehat{\mathcal{C}}^{\mathcal{P}}_{\xi}|=|\widehat{\mathcal{C}}^{\mathcal{Q}}_{\xi}|=|\widehat{\mathcal{P}}_{\xi}|=|\widehat{\mathcal{Q}}_{\xi}|$).
@@
As mentioned earlier, since the cylindrical feature is aligned with the local reference axes via PCA, the relative 3D rotation between ${\bm{v}}^{\bm{p}}_{3}$ (resp. ${\bm{v}}^{\bm{q}}_{3}$) and the unit $z$-axis, ${\bm{z}}=[0\;0\;1]^{\intercal}$, can be calculated using Rodrigues’ rotation formula [Mebius07arxiv-Derivation] as follows:
@@
where ${\bm{n}}^{{\bm{p}}}={\bm{v}}^{{\bm{p}}}_{3}\times{\bm{z}}$, $\theta^{{\bm{p}}}=\cos^{-1}\left({\bm{v}}^{{\bm{p}}}_{3}\cdot{\bm{z}}\right)$, and $\cdot$$_{\times}$ denotes the skew operator (resp. ${\bm{R}}^{\bm{q}}$). Thus, once the yaw rotation between the two patches ${\bm{R}}_{\text{yaw}}$ is determined, the full 3D rotation can be obtained as ${\bm{R}}=\left({\bm{R}}^{\bm{q}}\right)^{\intercal}{\bm{R}}_{\text{yaw}}{\bm{R}}^{\bm{p}}$.
As explained by Ao  et al. [Ao23CVPR-BUFFER], ${\bm{c}}^{{\bm{p}}}$ and ${\bm{c}}^{{\bm{q}}}$ follow discretized SO(2)-equivariant representation; thus, by finding the yaw rotation that maximizes circular cross-correlation between ${\bm{c}}^{{\bm{p}}}$ and ${\bm{c}}^{{\bm{q}}}$, we can estimate the relative SO(2) rotation ${\bm{R}}_{\text{yaw}}$. To this end, a 4D matching cost volume ${\bm{V}}\in\mathbb{R}^{H\times W\times W\times D}$ is constructed to represent the sector-wise differences between ${\bm{c}}^{{\bm{p}}}$ and ${\bm{c}}^{{\bm{q}}}$. Then, $\bm{V}$} is processed by a 3D cylindrical convolutional network (3DCCN) [Ao21cvpr-Spinnet], mapping $\bm{V}$} to a score vector $\bm{\beta}$} of size $W$.
@@
Cross-scale consensus maximization. Then, using per-pair $({\bm{R}},{\bm{t}})$ estimates from all scales, the 3D point pairs with the largest cardinality across scales should be selected as the final inlier correspondences $\mathcal{I}$, ensuring cross-scale consistency. To achieve this, we formulate the cross-scale inlier selection as consensus maximization problem [Sun22ral-TriVoC, Zhang24tpami-AcceleratingGloballyCM].
@@
$\text{s.t.}~~\|{\bm{R}}{\bm{p}}_{n}+{\bm{t}}-{\bm{q}}_{n}\|_{2}<\epsilon,~~\forall({\bm{p}}_{n},{\bm{q}}_{n})\in\mathcal{I}\subseteq\mathcal{D},$
@@
Finally, $\mathcal{I}$ is given as input to a solver, such as RANSAC [Fischler81] or TEASER++ [Yang20tro-teaser] to estimate $\hat{{\bm{R}}}$ and $\hat{{\bm{t}}}$. For a fair comparison with existing approaches, we use RANSAC.
@@ 4.4 Loss function and training
Loss functions. Unlike BUFFER, which was trained in four stages, our network follows a relatively simpler two-stage training process thanks to its detector-free nature. First, we train the feature discriminability of Mini-SpinNet descriptors using contrastive learning [Yew18eccv-3dfeatnet], followed by training $d$ in 7 to improve transformation estimation accuracy.
In particular, we employ the Huber loss [Zhang97ivc-Parameter] $\rho_{\text{Huber}}(\cdot)$ for training $d$ to remain robust to outliers[Barron19-generaladaptiverobustloss], while balancing sensitivity to small errors, which is formulated as follows:
@@
Patch distribution augmentation. Furthermore, we propose an inter-patch point distribution augmentation to allow Mini-SpinNet to experience a wider variety of patch distribution patterns. Specifically, we empirically sample the radius within $\frac{2}{3}r,\frac{4}{3}r]$ based on a uniform probability. As mentioned in Sec. 4.2, since $N_{\text{patch}}$ points within the radius are randomly selected as an input, a diverse set of patterns can be provided as $r$ varies.
@@ 5 Experiments
Datasets. As presented in Table A2, we designed our generalizability benchmark using eleven different datasets [Zeng17cvpr-3dmatch, Huang21cvpr-PREDATORRegistration, Yeshwanth23iccv-Scannet++, Qingqing22iros-TIERS, Sun20cvpr-WaymoDataset, Geiger13ijrr-KITTI, Pomerleau12ijrr-ETH, Jung23ijrr-HeLiPR, Tian23iros-KimeraMultiExperiments, Ramezani20iros-NewerCollege] to ensure balanced consideration of the following aspects: a) variation in environmental scales (i.e., indoor and outdoor environments), b) different scanning patterns with different sensor types, c) acquisition setups, and d) diversity of geographic and cultural environments as the data was collected across Europe, Asia, and the USA (i.e., Oxford, KAIST, and MIT campuses, respectively). More details can be found in Appendices B and C.
Training settings. Then, we only train the network on a single dataset, such as 3DMatch [Zeng17cvpr-3dmatch] or KITTI [Geiger13ijrr-KITTI]. Using the same hyperparameters of BUFFER [Ao23CVPR-BUFFER], we conducted a two-stage optimization (i.e., Mini-SpinNet is first trained, followed by training the 3DCCN) and we used Adam optimizer [Kingma2014arxiv-Adam] with a learning rate of 0.001, a weight decay of 1e-6, and a learning rate decay of 0.5. We used NVIDIA GeForce RTX 3090 with AMD EPYC 7763 64-Core.
Testing settings. In the case of the existing datasets [Zeng17cvpr-3dmatch, Huang21cvpr-PREDATORRegistration, Sun20cvpr-WaymoDataset, Geiger13ijrr-KITTI, Pomerleau12ijrr-ETH], we follow the conventional given pairs. The description of the newly employed datasets for evaluation can be found in Appendix B.
Evaluation Metrics. As a key metric, we use the success rate, which directly assesses the robustness of global registration [Lim24ijrr-Quatropp]. Specifically, a registration is deemed successful if the translation and rotation errors are within $\tau_{\text{trans}}$ and $\tau_{\text{rot}}$, respectively [Yew18eccv-3dfeatnet]. For successful cases, we evaluated the performance using relative translation error (RTE) and relative rotation error (RRE), which are defined as follows:
@@
Conven- tional FPFH [Rusu09icra-fast3Dkeypoints] + FGR [Zhou16eccv-FastGlobalRegistration] + [Uncaptioned image] 62.53 15.42 77.68 92.31 80.60 98.74 100.00 89.80 74.78 91.87 99.00 9.55
FPFH [Rusu09icra-fast3Dkeypoints] + Quatro [Lim22icra-Quatro] + [Uncaptioned image] 8.22 1.74 9.88 97.27 86.57 99.10 100.00 91.46 79.57 51.05 91.03 10.73
FPFH [Rusu09icra-fast3Dkeypoints] + TEASER++ [Yang20tro-teaser] + [Uncaptioned image] 52.00 13.25 66.15 97.22 73.13 98.92 100.00 89.20 71.30 93.69 99.34 10.00
Deep learning-based FCGF [Choi19iccv-FCGF] 88.18 40.09 72.90 88.69 55.96 0.00 0.00 0.00 0.00 54.98 0.00 15.00
@@
Ours 95.58 74.18 94.99 99.90 93.45 99.82 100.00 99.15 97.39 99.72 99.67 1.55
Table 1: Quantitative comparison of generalization performance in terms of success rate (%). Deep learning-based models were trained only on 3DMatch [Zeng17cvpr-3dmatch] and RANSAC was used with a maximum iteration of 50K. The icons represent oracle tuning ([Uncaptioned image]) for voxel size and radius, and scale alignment ([Uncaptioned image]) to normalize dataset scales to be similar to that of 3DMatch data (e.g., the scale of KITTI, which typically uses a voxel size of 0.3 m, is adjusted to match the scale of 3DMatch, where 0.025 m is commonly used, by dividing by $\frac{0.3}{0.025}$).
@@ 5.1 Analyses on the generalization
First, we demonstrate that existing methods struggle in achieving out-of-the-box generalization, leading to performance degradation due to the issues explained in Sec. 3. In this experiment, we mainly used renowned learning-based approaches: FCGF [Choi19iccv-FCGF], Predator [Huang21cvpr-PREDATORRegistration], GeoTransformer (GeoT for brevity) [Qin23tpami-GeoTransformer], BUFFER [Ao23CVPR-BUFFER], and PARENet [Yao24iccv-PARENet].
As shown in Table 1, models trained with small voxel sizes and search radii for indoor datasets exhibited substantial performance degradation in outdoor scenarios. In particular, as explained in Sec. 3.2.1, we observed that some approaches did not even work due to out-of-memory issues (i.e., N/A in Table 1) caused by an excessively large number of input points.
@@ 5.2 Performance comparison
Fig. 5: Relative translation error (RTE) and relative rotation error (RRE) of our approach to state-of-the-art methods, all trained on 3DMatch and tested on KITTI, with oracle tuning and scale alignment, corresponding to those in Table 1 under the + Refer to caption + Refer to caption setting. The **** annotations indicate measurements with a $p$-value $10^{-4}$ after a paired $t$-test.
@@
Conventional. G-ICP [Segal09rss-GeneralizedICP] 8.56 0.22 37.95
FPFH [Rusu09icra-fast3Dkeypoints] + FGR [Zhou16eccv-FastGlobalRegistration] 18.75 0.38 98.74
FPFH [Rusu09icra-fast3Dkeypoints] + Quatro [Lim22icra-Quatro] 18.56 0.93 99.10
FPFH [Rusu09icra-fast3Dkeypoints] + TEASER++ [Yang20tro-teaser] 15.35 0.68 98.92
KISS-Matcher [Lim25icra-KISSMatcher] 21.33 0.96 99.46
Learning-based 3DFeat-Net [Yew18eccv-3dfeatnet] 25.90 0.57 95.97
FCGF [Choi19iccv-FCGF] 6.47 0.23 98.92
DIP [Poiesi21icpr-DIP] 8.69 0.44 97.30
Predator [Huang21cvpr-PREDATORRegistration] 5.60 0.24 99.82
SpinNet [Ao21cvpr-Spinnet] 9.88 0.47 99.10
CoFiNet [Yu21nips-CofiNet] 8.20 0.41 99.82
D3Feat [Bai20cvpr-D3Feat] 11.00 0.24 99.82
GeDi [Poiesi22pami-GeDi] 7.55 0.33 99.82
GeoTransformer [Qin23tpami-GeoTransformer] 7.40 0.27 99.82
BUFFER [Ao23CVPR-BUFFER] 7.46 0.26 99.64
PARE-Net [Yao24iccv-PARENet] 4.90 0.23 99.82
Ours 7.74 0.27 99.82
Table 2: In-domain quantitative results in terms of relative translation error (RTE), relative rotation error (RRE), and success rate on KITTI [Yew18eccv-3dfeatnet].
@@ 5.3 Ablation study
Fig. 6: Effect of voxel size and search radius on success rates, where ★ indicates the best performance after tuning. (a) Impact of voxel size (with Refer to caption) on models trained on 3DMatch, evaluated on KITTI. (b) BUFFER-X vs. BUFFER across different radii, showing that optimal radius may vary within the same dataset.
Learning-based detector vs. Farthest point sampling. Interestingly, FPS rather showed better performance than the learning-based keypoint detector in BUFFER[Ao23CVPR-BUFFER], even in its training domain (i.e., in 3DMatch and 3DLoMatch). In particular, the performance gap sometimes becomes more pronounced in the out-of-domain scenes, demonstrating that robust cross-domain generalization can be achieved without the need for additional learning-based keypoint selection.
Fig. 7: Comparison of success rates between random sampling, learning-based keypoint detection in BUFFER [Ao23CVPR-BUFFER], and our farthest point sampling (FPS) strategy, showing that FPS performs comparably or even better across various datasets.